# Block-Attention for Efficient RAG

### v4

@@ Abstract
We introduce Block-Attention, an attention mechanism designed to address the increased inference latency and cost in Retrieval-Augmented Generation (RAG) scenarios. Traditional approaches often encode the entire context. Instead, Block-Attention divides retrieved documents into discrete blocks, with each block independently calculating key-value (KV) states except for the final block. In RAG scenarios, by defining each passage as a block, Block-Attention enables us to reuse the KV states of passages that have been seen before, thereby significantly reducing the latency and the computation overhead during inference. The implementation of Block-Attention involves block segmentation, position re-encoding, and fine-tuning the LLM to adapt to the Block-Attention mechanism. Experiments on four RAG benchmarks demonstrate that after block fine-tuning, the Block-Attention model achieves performance comparable to self-attention models (68.4% vs 67.9% on Llama3) or even superior performance (62.8% vs 59.6% on Mistral). Notably, Block-Attention significantly reduces the time to first token (TTFT) and floating point operations (FLOPs) to a very low level. It only takes 45 ms to output the first token for an input sequence with a total length of 32K. Compared to the self-attention models, the time consumption and corresponding FLOPs are reduced by 98.7% and 99.8%, respectively.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Li et al., 2022) is a crucial technology for mitigating knowledge hallucination in Large Language Models (LLMs). By utilizing retrieval technology, LLMs can seamlessly access passages stored in external databases, grounding their responses in the content of these passages. To the best of our understanding, RAG has emerged as the most effective method for infusing specific domain knowledge into LLMs in real-world scenarios.
@@
In this paper, we propose the Block-Attention method, which reduces the TTFT and computation FLOPs of RAG-LLMs to a level nearly equivalent to that of non-RAG LLMs, while fully maintaining the same accuracy level. As shown in Figure 1, the main idea of Block-Attention is to divide the entire input sequence into several blocks. Each block independently calculates its KV states through self-attention, without any attention to other blocks. Only the final block is able to attend other blocks, i.e., the user query is able to attend all the retrieved documents in the previous blocks. When utilizing Block-Attention in RAG scenarios, we may achieve substantial benefits by defining each passage as a single block and caching its KV states in the memory for further reuse.
@@
The implementation of Block-Attention can be easily achieved through the following steps: 1) Encode all blocks except the last one separately; 2) calculating the positional encoding for each based on its position in the input prompt; 3) Integrating these blocks to compute the KV states for the final block. However, the primary challenge in utilizing Block-Attention is that the LLMs have not been exposed to such an attention mechanism during their training, leading to difficulties in accurately interpreting the input prompt. In our preliminary experiments, we attempted to directly implement Block-Attention in the LLMs without updating any parameters. Unfortunately, this approach resulted in a substantial decrease in performance, with the average accuracy of Llama3-8B on four RAG benchmarks falling from 67.9% to 48.0%.
To address this challenge, we implemented a fine-tuning process for the LLMs to adapt to the Block-Attention mechanism. Our experiments demonstrated that, after approximately 100 - 1000 fine-tuning steps, the Block-Attention model achieved a full recovery of its original accuracy across both in-domain and out-domain scenarios, impressively increasing from 48.0% to 68.4%. This outcome underscores the Block-Attention LLMs’ capability to uphold inference accuracy while significantly enhancing inference efficiency in RAG scenarios.
We evaluate Block-Attention on four RAG benchmarks. Experimental results demonstrate that after fine-tuning, the average accuracy of the Block-Attention model on the benchmarks can remain comparable to (68.4% vs 67.9% on Llama3) or even slightly exceed (62.3% vs 59.6% on Mistral) the performance of self-attention models. In terms of efficiency, we counted the TTFT and FLOPs to the first token (FLOPs-TFT) of the block-attention model when the length of user input is 50 and the total length of the input sequence gradually increases. We found that the longer the total length, the more obvious the improvement of block-attention on efficiency. When the length of the input sequence reaches 32K, the TTFT and FLOPs-TFT of the block-attention model are 1.3% and 0.2% of that of the self-attention model, respectively. [^1] Given that the KV cache is already a mature and low-cost technology (Qin et al., 2024; Lan et al., 2023; Lee et al., 2021), in this paper we do not take the cost of KV cache into account.
@@ 2.1 Main Idea
In the scenario of RAG, a given passage frequently recurs in the retrieval results for numerous queries (Qin et al., 2024), making it a natural consideration to explore the possibility of reusing its KV states. Unfortunately, the auto-regressive nature of the encoding approach, as previously outlined, imposes constraints that render the KV states nearly non-reusable across distinct queries. Our research is dedicated to examining a novel attention mechanism, Block-Attention, that designed to necessitate the re-computation of only the altered text blocks between two input sequences, thereby achieving an outcome equivalent to that of full sequence re-encoding.
As illustrated in Figure 1, the essence of Block-Attention is to divide the input sequence $\mathcal{S}$ into several independent blocks. Each block autonomously calculates its KV states through self-attention, without considering other blocks. The final block, however, has the unique capability to integrate information from preceding blocks. A primary advantage of this method is the modular independence it provides: when a block $b_{i}$ is updated to $b^{\prime}_{i}$, re-encoding only the KV states of the affected block $k_{b^{\prime}_{i}}$, $v_{b^{\prime}_{i}}$, and those of the final block $k_{b_{k}}$, $v_{b_{k}}$, is sufficient to obtain the updated sequence’s KV states.
To develop a Block-Attention LLM capable of precise inference, we must tackle three key challenges:
  • How do we segment blocks?
  • How should the positional encoding be calculated for each block?
  • How can the LLM be adapted to the Block-Attention mechanism?
@@ 2.2 Block Segmentation
The primary principle of block division is to segment semantically independent parts of the prompt into separate blocks. In RAG scenarios, since the retrieved passages are originally mutually independent, it is natural to divide them into different blocks. Therefore, let’s go back to the left part of Figure 1, where we allocate each passage to a single block and designate the user’s query as the final block. This principle extends to other scenarios as well. For example, in the context of code generation tasks, a function may be treated as one block; in multi-turn dialogues, each turn could be segmented into an individual block. In this paper, our primary focus is on the application of Block-Attention in RAG, with the exploration of other scenarios reserved for future research.
@@ 2.4 Block Fine-tune
Due to the LLM’s reliance on full self-attention during the training phase, a direct switch to Block-Attention during inference might result in a significant discrepancy between the training and inference states. Our preliminary findings indicate that introducing Block-Attention without subsequent fine-tuning could precipitate a substantial decrease in performance, with the average accuracy dropping significantly from 67.9% to 48.0%. Adapting the LLM to Block-Attention through fine-tuning, which we refer to as ”block fine-tune,” is quite straightforward. The only difference from the standard SFT process is the modification of the traditional lower-triangular attention mask matrix to the attention mask matrix depicted in the right part of Figure 1. With this masking matrix, tokens in all blocks except the last are restricted to attending only to information within their own block, thereby ensuring consistency between training and inference.
Figure 2: The Inference Pipeline of Block-Attention Model
@@ 3 Experiments
After the above analysis, there exists two concerns about the Block-Attention method: 1) Can the Block-Attention model achieve the same accuracy as self-attention in RAG scenarios? 2) How much can the Block-Attention mechanism improve the efficiency? The following experimental results will reveal the answers to these two questions. In Sections 3.5, we analyze the accuracy of Block-Attention models. Meanwhile, in Section 3.6, we demonstrate the efficiency of Block-Attention.
@@ 3.1 Datasets
We construct the training datasets by randomly sampling 20,000 instances from these TriviaQA (TQA) (Joshi et al., 2017) and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020) for fine-tuning models. Each training sample consists of (1) a question, (2) 10 passages retrieved from these two datasets using the Contriver toolkit [^2] https://github.com/facebookresearch/contriever, which identifies the 10 most relevant passages, and (3) an answer generated by GPT-4 based on the retrieved passages. The reason to use the GPT-4 answer instead of the ground-truth answers is because the answer might not be present in our retrieved passages. This discrepancy could lead the model to overlook the content of the retrieved passages and generate outputs directly.
@@
We evaluate the performance of our proposed Block-Attention mechanism and baseline models on four widely-used RAG benchmarks: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), HotpotQA (HQA) (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). Since the training samples include instances from TQA and 2Wiki, their test sets are considered in-domain, whereas HQA and NQ are treated as out-of-domain test sets. Following  Kandpal et al. (2023) and  Liu et al. (2024), we use accuracy as our primary evaluation metric, judging whether any correct answers appear in the predicted output. To mitigate biases arising from output length, we set a maximum token limit of 200 for the output sequences.
@@ 3.3 Base Model & Baselines
We implement the Block-Attention mechanism on two base language models: Llama3-8B-base [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B and Mistral-7B-v3.0444https://huggingface.co/mistralai/Mistral-7B-v0.3. After applying block fine-tuning, these models are denoted as Llama3-block-ft and Mistral-block-ft. For comparison, we implemented three baselines for each of these two base language models.
  • Strong baselines: Llama3-vanilla-sft and Mistral-vanilla-sft are established by fine-tuning the base language models on our constructed training dataset. These models serve as the upper bound performance for the Block-Attention models. We aim for the Block-Attention model’s performance to approach these as closely as possible.
@@
  • Weak baselines: Llama3-block-w/o-ft and Mistral-block-w/o-ft. These models transition the attention mechanism of Llama3-vanilla-sft and Mistral-vanilla-sft to a Block-Attention mechanism without any block fine-tuning. The outcomes of these models represent the lower bounds for the Block-Attention model’s effectiveness, given that they have not undergone any adaptation to Block-Attention during the training phase.
@@
  • w/o-pos: Llama3-block-ft-w/o-pos and Mistral-block-ft-w/o-pos are baselines that also fine-tuned to adapt the Block-Attention mechanism, while no additional position re-encoding operations described in Section 2.3 are conducted. This group of models will be used to evaluate the effectiveness of the proposed position re-encoding process.
Please note that we have not utilized instruction-tuning models as our base language model, such as Llama3-instruct. The reason is our inability to entirely re-implement the fine-tuning processes of these models, which precludes a direct comparison between self-attention and Block-Attention methodologies under identical experimental conditions. Instead, we perform vanilla fine-tuning and block fine-tuning on the same base models using the same dataset, ensuring a fair comparison.
Additionally, our idea has some similarities with PCW  (Ratner et al., 2023b) and Prompt Cache (Gim et al., 2024). However, the focus of PCW is on extending the context window rather than efficient inference. When we directly applied their attention mechanism in the RAG scenario, their accuracy was even lower than the weak baselines, so we did not choose it as a baseline. We also re-implemented Prompt Cache. Then, we had an interesting finding: their results are exactly the same as the two weak baselines: Llama3-block-w/o-ft and Mistral-block-w/o-ft. Therefore, for the results of Prompt Cache, please directly refer to the result of these two models. [^5] Please note that in addition to accuracy, Prompt Cache also has serious efficiency issues when applied in the RAG field: once the rank of a passage in the input sequence changes, its KV cache cannot be reused.
3.4 Setup
@@
All experiments are done with 8 NVIDIA A800 GPUs, we set the learning rate $\alpha=2\times 10^{-5}$, batch size $b=64$ and the number of epochs $n=1$ for all models. We use a linear learning rate warmup for the first 20 steps. Its training is in bfloat16 format, using DeepSpeed ZeRO Stage-2666https://github.com/microsoft/DeepSpeed and Flash-Attention V2 (Dao et al., 2022). The Flash-Attention toolkit is used for efficient inference for our models and baselines.
Models TQA 2wiki NQ HQA
In-Dom.
Avg.
Out-Dom.
Avg.
Mix
Avg.
Llama3-vanilla-sft 72.8 73.4 55.7 69.5 73.9 62.0 67.9
Llama3-block-ft 73.0 73.3 56.2 68.5 73.6 63.2 68.4
Llama3-block-ft-w/o-pos 70.8 66.8 52.9 61.4 69.3 58.5 63.9
Llama3-block-w/o-ft 62.7 48.1 43.3 39.6 54.7 41.4 48.0
Mistral-vanilla-sft 66.6 69.5 46.0 56.4 68.3 50.9 59.6
Mistral-block-ft 69.9 70.5 50.8 57.9 70.2 54.4 62.3
Mistral-block-ft-w/o-pos 67.8 67.4 47.5 52.4 67.3 50.4 58.8
Mistral-block-w/o-ft 51.9 44.6 32.1 27.8 47.7 29.7 38.7
Table 1: Accuracy of different models on four benchmarks. Please note that the “average” column does not depict a simple arithmetic mean of the preceding columns, as the number of samples contained in each test set varies.

### v5

# Block-attention for Efficient Prefilling

@@ Abstract
We introduce Block-attention, an attention mechanism designed to address the increased inference latency and cost in Retrieval-Augmented Generation (RAG) scenarios. Traditional approaches often encode the entire context in an auto-regressive manner. Instead, Block-attention divides retrieved documents into discrete blocks, with each block independently calculating key-value (KV) states except for the final block. In RAG scenarios, by defining each passage as a block, Block-attention enables us to reuse the KV states of passages that have been seen before, thereby significantly reducing the latency and the computation overhead during inference. The implementation of Block-attention involves block segmentation, position re-encoding, and fine-tuning the LLM to adapt to the Block-attention mechanism. Experiments on 11 diverse benchmarks, including RAG, ICL, and general domains, demonstrate that after block fine-tuning, the Block-attention model not only achieves performance comparable to that of full-attention models, but can also seamlessly switch between the block and full attention modes without any performance loss. Notably, Block-attention significantly reduces the time to first token (TTFT) and floating point operations (FLOPs) to a very low level. It only takes 45 ms to output the first token for an input sequence with a total length of 32K. Compared to the full-attention models, the TTFT and corresponding FLOPs are reduced by 98.7% and 99.8%, respectively. Additionally, in Appendix A, we elaborate on how Block-attention is applied in Game AI scenario and the substantial potential benefits it entails. We strongly suggest researchers in the gaming field not to overlook this section.  [^1] Codes, datasets and model weights have been publicly available at https://github.com/TemporaryLoRA/Block-attention.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Li et al., 2022; Lan et al., 2023) is a crucial technology for mitigating knowledge hallucination in Large Language Models (LLMs). By utilizing retrieval technology, LLMs can seamlessly access passages stored in external databases, grounding their responses in the content of these passages. To the best of our understanding, RAG has emerged as the most effective method for infusing specific domain knowledge into LLMs in real-world scenarios.
@@
In this paper, we propose the Block-attention method, which reduces the TTFT and computation FLOPs of RAG-LLMs to a level nearly equivalent to that of non-RAG LLMs, while fully maintaining the same accuracy level. As shown in Figure 1, the main idea of Block-attention is to divide the entire input sequence into several blocks. Each block independently calculates its KV states through full-attention, without any attention to other blocks. Only the final block is able to attend other blocks, i.e., the user query is able to attend all the retrieved documents in the previous blocks. When utilizing Block-attention in RAG scenarios, we may achieve substantial benefits by defining each passage as a single block and caching its KV states in the memory for further reuse.
@@
The implementation of Block-attention can be easily achieved through the following steps: 1) Encode all blocks except the last one separately; 2) calculating the positional encoding for each based on its position in the input prompt; 3) Integrating these blocks to compute the KV states for the final block. However, the primary challenge in utilizing Block-attention is that the LLMs have not been exposed to such an attention mechanism during their training, leading to difficulties in accurately interpreting the input prompt [^2] Although some studies have proposed training-free frameworks (Gim et al., 2024; Merth et al., 2024) utilizing parallel context encoding and KV caching to achieve passage-level KV cache reuse, they encounter significant degeneration issues in practical applications. Our experimental results also demonstrate that the results of these methods are far from satisfactory.. In our preliminary experiments, we attempted to directly implement Block-attention in the LLMs without updating any parameters. Unfortunately, this approach resulted in a substantial decrease in performance, with the average accuracy of Llama-3.1-Tulu-3-8B-SFT on four RAG benchmarks falling from 66.1% to 49.9%.  [^3] We also proposed some training-free solutions to mitigate this huge gap. If someone is interested in switching to Block-attention in an on-the-fly manner, please refer to Zhang et al. (2024) for the detail.
To address this challenge, we implemented a fine-tuning process for the LLMs to adapt to the Block-attention mechanism. Our experiments demonstrated that, after approximately 500-1000 fine-tuning steps, the Block-attention model achieved a full recovery of its original accuracy across all scenarios, impressively increasing from 49.9% to 70.0%. This outcome underscores the Block-attention LLMs’ capability to uphold inference accuracy while significantly enhancing inference efficiency in RAG scenarios.
We conduct comprehensive evaluations of the Block-attention mechanism across 11 diverse benchmarks, including RAG, ICL(In-Context Learning), and general domains. Experimental results demonstrate that after fine-tuning, the average accuracy of the Block-attention model on the benchmarks can remain comparable to Llama-3.1-Tulu-3-8B-SFT. In terms of efficiency, we counted the TTFT and FLOPs to the first token (FLOPs-TFT) of the Block-attention model when the length of user input is 50 and the total length of the input sequence gradually increases. We found that the longer the total length, the more obvious the improvement of Block-attention on efficiency. When the length of the input sequence reaches 32K, the TTFT and FLOPs-TFT of the Block-attention model are 1.3% and 0.2% of that of the full-attention model, respectively. [^4] Given that the KV cache is already a mature and low-cost technology (Qin et al., 2024; Lee et al., 2021), in this paper we do not take the cost of KV cache into account.
@@ 2.1 Main Idea
Our research focuses on exploring a novel attention mechanism named Block-attention. This mechanism is designed in such a way that it only requires the re - computation of the text blocks that have been altered between two input sequences. As a result, it can achieve an outcome that is equivalent to that obtained from fully re - encoding the entire sequence.
As illustrated in Figure 1, the essence of Block-attention is to divide the input sequence $\mathcal{S}$ into several independent blocks. Each block autonomously calculates its KV states through self-attention, without considering other blocks. The final block, however, has the unique capability to integrate information from preceding blocks. A primary advantage of this method is the modular independence it provides: when a block $b_{i}$ is updated to $b^{\prime}_{i}$, re-encoding only the KV states of the affected block $k_{b^{\prime}_{i}}$, $v_{b^{\prime}_{i}}$, and those of the final block $k_{b_{k}}$, $v_{b_{k}}$, is sufficient to obtain the updated KV states.
To develop a Block-attention LLM capable of precise inference, we must tackle three challenges:
1) How do we segment blocks?
2) How should the positional encoding be calculated for each block?
3) How can the LLM be adapted to the Block-attention mechanism?
@@ 2.2 Block Segmentation
The primary principle of block division is to segment semantically independent parts of the prompt into separate blocks. In RAG scenarios, since the retrieved passages are originally mutually independent, it is natural to divide them into different blocks. Therefore, let’s go back to the left part of Figure 1, where we allocate each passage to a single block and designate the user’s query as the final block. This principle extends to other scenarios as well. For example, in the context of code generation tasks, a function may be treated as one block; in multi-turn dialogues, each turn could be segmented into an individual block; while in In-context Learning (ICL), each demonstration naturally forms a self-contained block (we will validate the efficacy of Block-attention in ICL scenarios in our experiments). In this paper, our primary focus is on the application of Block-attention in RAG, with the exploration of other scenarios reserved for future research.
@@ 2.3 Position Re-Encoding
Figure 2: The Inference Pipeline of Block-attention Model
@@ 2.4 Block Fine-tune
Due to the LLM’s reliance on full-attention during the training phase, a direct switch to Block-attention during inference might result in a significant discrepancy between the training and inference states. Our preliminary findings indicate that introducing Block-attention without subsequent fine-tuning could precipitate a substantial decrease in performance, with the average accuracy dropping significantly from 67.9% to 48.0%. Adapting the LLM to Block-attention through fine-tuning, which we refer to as ”block fine-tune,” is quite straightforward. The only difference from the standard SFT process is the modification of the traditional lower-triangular attention mask matrix to the attention mask matrix depicted in the right part of Figure 1. With this masking matrix, tokens in all blocks except the last are restricted to attending only to information within their own block, thereby ensuring consistency between training and inference.
@@ 3 Experiments
After the above analysis, there exist three concerns about the Block-attention method: 1) Can the Block-attention model achieve the same accuracy as full-attention in multi-block scenarios such as RAG and ICL? 2) Can the Block-attention model still adapt to the full-attention mechanism? 3) How much can the Block-attention mechanism improve the efficiency? The following experimental results will reveal the answers to these three questions. In Sections 3.5, we explored the answers to the first two questions and analyzed the accuracy of Block-attention models in 11 diverse tasks. Meanwhile, in Section 3.6, we demonstrate the efficiency of Block-attention in RAG scenarios.
@@ 3.1 Datasets
[…]
Another part of our training dataset is RAG samples. We randomly sample 20,000 instances from TriviaQA(TQA) (Joshi et al., 2017) and 2WikiMultiHopQA(2Wiki) (Ho et al., 2020) for fine-tuning models. Each training sample consists of (1) a question, (2) 10 passages retrieved from these two datasets using the Contriever toolkit [^6] https://github.com/facebookresearch/contriever, which identifies the 10 most relevant passages, and (3) an answer generated by Llama3.3-70B-Instruct based on the retrieved passages. The reason for using the Llama3 answer instead of the ground-truth answers is that the answer might not be present in our retrieved passages. This discrepancy could lead the model to overlook the content of the retrieved passages and generate outputs directly.
In the experiment, to maintain the full-attention ability of the model, we train the Block-attention model using both full-attention and Block-attention mechanism simultaneously. [^7] The details about our processed datasets can be found at https://github.com/TemporaryLoRA/Block-Attention. In other words, all samples in the training set will be trained in both ways. For the full-attention baselines, only the full-attention method is used for training.
@@
We evaluate the performance of our proposed Block-attention mechanism and baseline models on four widely-used RAG benchmarks: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), HotpotQA (HQA) (Yang et al., 2018), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), and NarrtiveQA (NQA) (Kočiský et al., 2017). Following  Kandpal et al. (2023) and  Liu et al. (2024), we use accuracy as our primary evaluation metric, judging whether any correct answers appear in the predicted output. To mitigate biases arising from output length, we set a maximum token limit of 200 for the output sequences.
In addition, we also evaluated the performance of the Block-attention model and the full-attention model in seven benchmarks of the general domain: MMLU (Hendrycks et al., 2021a), BigBenchHard (BBH) (Suzgun et al., 2022), DROP (Dua et al., 2019), MATH (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and IFEval (Zhou et al., 2023).  [^8] Since we do not have access to the OpenAI API, we are unable to conduct evaluations on datasets that require subjective assessment. Please note that, among them, BBH, DROP, GSM8K, and MATH are ICL tasks with several independent in-context examples. We will divide each example into a separate block. For zero-shot datasets like MMLU, HumanEval, and IFEval, the block attention model will directly switch to the full-attention mode.
@@ 3.3 Base Model & Baselines
We implement the Block-attention mechanism on Llama-3.1-Tulu-3-8B-SFT (denoted as Tulu3-SFT) [^9] https://huggingface.co/allenai/Llama-3.1-Tulu-3-8B-SFT. The reason why we choose this model is that it is a model that can be reproduced (They released their SFT data). By comparing different fine-tuning methods with the same data, we can easily demonstrate the effectiveness of Block-attention in different scenarios. After applying block fine-tuning, the model is denoted as Tulu3-block-ft. For comparison, we also implemented four baselines and three ablated models.
  • Tulu3-SFT: Our base model, which also serves as the performance ceiling for Block-attention models in general tasks. We aim to make the Block-attention model’s general performance approach this benchmark as closely as possible.
  • Tulu3-RAG: Since Tulu3-SFT was not trained on RAG data, it demonstrates suboptimal performance in RAG tasks. To ensure fair comparison, we conducted supervised fine-tuning of Tulu3-SFT using our training data with full-attention mechanism. This model establishes the performance ceiling for Block-attention models in RAG scenarios. We similarly aim to have the Block-attention model’s RAG performance approach this upper bound as closely as possible.
  • Tulu3-RAG-promptCache: A baseline model that applies the PromptCache (Gim et al., 2024) to the Tulu3-RAG model. This enables training-free reuse of attention states across different prompts of large language models (LLMs).
  • Tulu3-RAG-Superposition: Serving as a baseline model, it uses the Superposition method on the Tulu3-RAG model (Merth et al., 2024). This method allows the LLM to process input documents along parallel prompt paths. This parallel processing mechanism not only enhances the model’s processing speed but also optimizes resource utilization by eliminating unnecessary computations.
  • Tulu3-block-ft-full. An ablated model that switches the Tulu3-block-ft to the full-attention mode. We want to observe through the performance of this ablated model whether the Block-attention and the full-attention can coexist in the same model.
@@
  • Tulu3-block-w/o-ft. An ablated model that transitions the attention mechanism of Tulu3-RAG to Block-attention mechanism without any block fine-tuning. The outcomes of this model represent the lower bounds for the Block-attention model’s effectiveness, given that it has not undergone any adaptation to Block-attention during the training phase.
@@
  • Tulu3-block-w/o-pos. Another ablated model that is also fine-tuned to adapt the Block-attention mechanism, while no additional position re-encoding operations described in Section 2.3 are conducted. This model will be used to evaluate the effectiveness of the proposed position re-encoding process.
3.4 Experimental Setup
Training and Inference
@@ 3.4 Experimental Setup
All experiments are conducted using 8 NVIDIA H20 GPUs with following hyper-parameters: (1) learning rate $\alpha=2\times 10^{-5}$; (2) batch size $b=64$; (3) epochs $n=1$; and (4) 20 warmup steps. The DeepSpeed [^10] https://github.com/microsoft/DeepSpeed and Flash-Attention (Dao et al., 2022) toolkits are utilized to accelerate our training procedure using bfloat16 format. Additionally, Flash-Attention is utilized for efficient inference of our fine-tuned models and baselines.
Evaluation
OpenCompass toolkit (Contributors, 2023) is used to evaluate the performance of models under general benchmarks and RAG benchmarks.
Models 2wiki HQA NQ TQA
Tulu3-SFT 62.0 68.4 58.6 75.7
Tulu3-RAG 73.2 74.8 61.5 75.8
Tulu3-RAG-Superposition 30.1 32.3 35.9 58.9
Tulu3-RAG-promptCache 32.4 31.6 44.4 61.8
Tulu3-block-ft 72.2 72.3 60.4 75.1
Tulu3-block-ft-full 73.6 75.2 62.2 76.2
Tulu3-block-ft-w/o-pos 68.9 69.9 59.2 74.4
Tulu3-block-w/o-ft 42.9 42.1 48.3 66.5
Table 1: Accuracy of different models on four RAG benchmarks.
Task Type General ICL
dataset IFEval HumanEval MMLU GMS8K MATH BBH DROP
setup 0-shot 0-shot 0-shot 4-shot 4-shot 3-shot 3-shot
Tulu3-SFT 68.5 58.5 63.7 75.5 29.2 68.5 9.4
Tulu3-RAG 68.3 65.2 63.6 75.6 28.6 68.5 10.4
Tulu3-block-ft 70.0 59.1 63.0 75.7 28.8 65.3 14.4
Table 2: Accuracy of different models on seven general benchmarks. For the first three zero-shot benchmarks, the Block-attention will fall back to full-attention. For the subsequent four ICL datasets with few-shot examples, each sample will be divided into an independent block. Therefore, for a k-shot sample, it will be divided into k+1 blocks.